OpenStudy (anonymous):

domain and range F(X)=-x^2+2X+3

OpenStudy (anonymous):

\[F \left( x \right)=x^2+2x+3=x^2+2x+1+2=\left( x+1 \right)^2+2\] domain= all real values \[range=[2,\infty ) \]

OpenStudy (anonymous):

@surjithayer, \(x^2\) is negative

OpenStudy (anonymous):

\[F \left( x \right)=-x^2+2x+3=-\left( x^2-2x+1-1 \right)+3=-\left( x-1 \right)^2+4\] domain is same. \[Range=(-\infty ,4]\]